# GiVE: Guiding Visual Encoder to Perceive Overlooked Information

### v1

@@ Introduction
AG-CLIP 62M CLIP-ViT B/32 -- 29.1 29.8 50.7 75.6 56.2 77.3
Table 3: The effect of semantic instructions on secondary coding methods. AG-CLIP is CLIP-ViT with GiVE for attention guidance. “Inst.” indicates whether the model receives additional instructions. “Img Enc.” specifies the image encoder used by the model, and “Type” denotes the type of image encoder. “GiVE” shows whether our GiVE method is applied to the image encoder, and “Post Proc.” identifies the type of post-processor used for the secondary encoding of the image embeddings.
Experiments
Experimental Setup
Datasets
We evaluate the effectiveness of GiVE using both the LVIS dataset (Gupta, Dollár, and Girshick 2019) and MOInst dataset, each annotating multiple objects per image. The comprehensive LVIS dataset, referred to as LVISa, comprises 1,203 categories, from which, we derive a subset, denoted as LVISf, with 405 “frequent” categories. In contrast, the MOInst dataset used for training contains 264 categories, with category label overlaps of 7.7% and 17.6% with the two LVIS datasets, respectively. We also conduct experiments on the COCO (Lin et al. 2014), SUN397 (Xiao et al. 2010), and ImageNet (Gao et al. 2023) datasets.
Baselines
We evaluate the gains brought by integrating the GiVE with several representative ViT baselines, including CLIP (Radford et al. 2021), GroupViT (Xu et al. 2022), SigLIP (Zhai et al. 2023), and OwlViT (Minderer et al. 2022). GroupViT has two versions trained on different datasets, and OwlViT is a visual encoder designed for object detection. Additionally, we apply the GiVE to BLIP-2 (Li et al. 2023) and InstructBLIP (Dai et al. 2023) to explore its impact within a larger encoding framework. Throughout this paper, unless explicitly stated otherwise, the experiments with GiVE are conducted on the CLIP platform, with the unmarked CLIP model referring to CLIP-ViT-B/32. Furthermore, a comparison is presented between CLIP using GiVE and the more recent MetaCLIP (Xu et al. 2024).
Evaluation Metrics
Inspired by prior works (Li et al. 2023; Yao et al. 2022), we conduct quantitative evaluations on image classification and image-text retrieval tasks. Scores are calculated solely from vector similarity, without further optimization, to accurately reflect the effective information content of the extracted features. To ensure that the scores reflect the visual information extraction capabilities of visual encoders, we mitigate the potential interference by quantifying the influence of textual data when evaluating methods that use instruction text, such as InstructBLIP (Dai et al. 2023). For detailed calculations and implementation specifics, see the Appendix.
Figure 4: Visual encoder attention visualization. The images are divided into three groups, each containing the raw image and its attention visualization. The first row shows the raw image and the attention of CLIP-ViT. The second row displays attention maps of CLIP integrated with the AG-Adapter under various instructions.
Loss Fuse Inst. MOInst LVISf LVISa
@@
-OITC early+late 0.1 0.1 0.3 50.0 0.1 50.0
@@
all early+late (dense) 29.1 29.8 50.7 75.6 56.2 77.3
Table 4: Ablation study. The “fuse” represents the layer where the AG-Adapter is inserted. The “early” indicates feature fusion in only the first half of the layers near the input, while the “late” indicates fusion in the latter half. The “sparse” refers to fusion in alternating layers, whereas the “dense” means fusion in all layers.
Performance Evaluation
Image Classification
We conduct zero-shot image classification on the “frequent” subset and the full LVIS test set. Table 1 presents the results on five ViT baselines. Key observations from these experiments are as follows:
@@
Please see Table 11 in the Appendix for a comparison with MetaCLIP on more datasets.
Image-Text Retrieval
Image-text retrieval includes two subtasks: image-to-text retrieval and text-to-image retrieval. We evaluate the models on the MOInst dataset. Since the AG-Adapter is trained on this dataset, we compare the results under the fine-tuned setting, as recorded in Table 2.
As shown in Table 2, GiVE outperforms all baseline methods while utilizing fewer trainable parameters. And the informativeness gap between instruction and caption, illustrated in Figure 5 of Appendix, indicates that simple instruction alone is insufficient for the model to achieve such a high hit rate. Together, GiVE effectively extracts target visual features. Additionally, the evaluation results for “Instruct.” further support the credibility of our experimental findings.
Comparison with Secondary Coding Methods
The BLIP-2 and InstructBLIP methods utilize a Q-Former to re-encode image features after the visual encoder, with InstructBLIP incorporating text instructions into the Q-Former. We replace the visual encoders of these two methods with AG-CLIP-ViT-B/32 and then fine-tune them on MOInst following their respective strategies. The results are shown in Table 3. From the table, it can be observed that:
@@
Ablation Studies
In this section, we conduct an ablation study of the GiVE from three perspectives: loss, fusion layers, and abstract semantic instructions. Specifically, for instructions, we eliminate abstract semantics by replacing short object prompts with detailed descriptions. Table 4 presents the experimental results, with the gray row representing the original CLIP-ViT and the yellow row representing our AG-CLIP-ViT. Observations from the results are as follows:
@@
Attention Visualization
In this section, we use EigenCAM (Muhammad and Yeasin 2020) to visualize the attention of CLIP-ViT with and without the integration of the GiVE on three example images in Figure 4. It is evident that GiVE aids the original ViT in encoding images more effectively in the following ways:
  • Flexibility for a single image. The dynamically changing attention areas in each image indicate that the ViT no longer encodes fixed image features but can extract visual features tailored to the query.
  • Effectiveness of semantic control over attention. The three attention distributions for the left and middle images show that even objects initially overlooked by the ViT can be brought into focus through instructions.
  • Completeness of object retrieval. As shown in the right image with the “shoe” instruction, all shoes are attended to, regardless of whether they are in a box or separated by a human body. GiVE enables the ViT to locate more comprehensive specific objects.
For additional results, see Figure 7 in the Appendix.
Conclusion
@@
Appendix A Appendix
The appendix covers several key topics: first, it provides an example of the inference pipeline to illustrate the GiVE workflow. Next, it presents statistics on the types of image encoders used by MLLMs. Following this, it describes the proposed MOInst dataset and its construction process. The appendix then details the experimental setup and includes additional experiments, such as parameter sensitivity and visualization illustrations. Finally, it demonstrates the impact of visual attention on MLLMs and the improvements achieved with our GiVE approach.
Prompt Engineering and Pipeline
GiVE aims to provide text guided visual coding scheme. Text is a medium that is readily comprehensible and straightforward for both humans and models. Therefore, we chose to utilize text, which introduces the challenge of avoiding information leakage. To mitigate this issue, we provide only minimal information to the visual encoder, specifically words that serve as object labels. This level of information is sufficient to prevent the visual encoder from missing key features.
However, text encoders are typically trained on complete sentences describing images rather than single words. To address this discrepancy, object indicator words are incorporated into the template “A photo of {object}” during both training and inference. Figure 5 illustrates the pipeline of the model as an example of image-text retrieval.
Figure 5: Pipeline for image-text retrieval. The frozen text encoder synthesizes retrieval instructions through constructed statements with target objects. The trained image encoder generates corresponding image embeddings for different instructions.
MLLM and Image Encoder Summary
An investigation was conducted into 30 common MLLMs from the past two years, with a view to establish which image encoders they employ and the pre-training tasks associated with these encoders. The findings are set out in Table 5. It is evident that the majority of researchers prefer pre-aligned encoders such as CLIP (Radford et al. 2021) when constructing MLLM models. Even the few models that employ image reconstruction encoders for specific purposes, such as expanding modalities (Yang et al. 2023) or freezing LLMs (Liu, Yan, and Abbeel 2023; Yu et al. 2023), exhibit diminished capacity for multimodal understanding in comparison to models like LLaVA (Liu et al. 2023b), whose encoders undergo training with contrastive learning.
Model Image Encoder Pre-training Task
BLIP-2 (2023) CLIP (2021) /EVA-CLIP (2023) contrast
InstructBLIP (2023) CLIP (2021) /EVA-CLIP (2023) contrast
X-InstructBLIP (2023) EVA-CLIP (2023) contrast
LLaVA (2023b) CLIP (2021) contrast
LLaVA-1.5 (2023a) CLIP (2021) contrast
LLaVA-NeXT (2024) CLIP (2021) contrast
BLIVA (2024) EVA-CLIP (2023) contrast
MiniGPT-4 (2023) EVA-CLIP (2023) contrast
MiniGPT-5 (2023) EVA-CLIP (2023) contrast
MiniGPT-v2 (2023a) EVA-CLIP (2023) contrast
mPLUG-Owl (2023) CLIP (2021) contrast
mPLUG-Owl2 (2024) CLIP (2021) contrast
ShareGPT4V (2023c) CLIP (2021) contrast
PandaGPT (2023) ImageBind (2023) contrast
Video-LLaMA (2023) EVA-CLIP (2023) contrast
Video-ChatGPT (2023) CLIP (2021) contrast
Shikra (2023b) CLIP (2021) contrast
VisualGLM-6B (2023) EVA-CLIP (2023) contrast
CogVLM (2023) EVA-CLIP (2023) contrast
Mini-Gemini (2024) CLIP (2021); ConvNeXt (2022) contrast
BuboGPT (2023) CLIP (2021) /EVA-CLIP (2023) contrast
ChatSpot (2023) CLIP (2021) contrast
Qwen-VL (2023) OpenCLIP (2023) contrast
NExT-GPT (2023) ImageBind (2023) contrast
DRESS (2023d) EVA-CLIP (2023) contrast
Cobra (2024) SigLIP (2023); DINOv2 (2023) contrast
CoDi-2 (2024) ImageBind (2023) contrast
LQAE (2023) LQAE (2023) reconstruction
SPAE (2023) SPAE (2023) reconstruction
TEAL (2023) BEiT-V2 (2022) reconstruction
Table 5: A summary of common MLLMs and the image encoders they use. This table presents the image encoder pre-training tasks. Most researchers prefer using pre-aligned encoders when building MLLM models.
Multi-Object Instruction Dataset
Our MOInst dataset is divided into training and testing sets, and the specific statistics are shown in Table 6. The annotation files of Multi-Object Instruction (MOInst) encompass the following fields: “image_id” indicates the unique identifier of a particular image, “file_name” serves as the image file name, “image_path” specifies the path to the image file, “caption” provides a textual description of an object in the image, “category_id” corresponds to the category identifier of the caption, and “object_id” uniquely identifies a particular category within a specific image. The key information for an image in the dataset is shown in Table 7 as an example.
Parts # Imgs # Caps # Cats # Avg. Cats/Image
Train 80536 240183 264 2.4
Test 1000 4195 264 3.2
Table 6: Statistics of the MOInst dataset. Parts represents the different partitions of the dataset, # Imgs indicates the total number of images in each partition, # Caps refers to the total number of captions, # Cats denotes the total number of distinct categories, and # Avg. Cats/Image is the average number of categories labeled per image. Multiple captions of an image may share the same category label.
Construction
We refer to the pipeline of (Chen et al. 2023e) to construct our dataset. As shown in Figure 6, the pipeline consists of two stages. The first stage aims at generating category labels for the objects described by captions, while the second stage improves the quality of the labels.
Figure 6: Pipeline for dataset construction. The dataset construction includes both generation and processing stages. It is created in a semi-automated manner with the help of existing models and minimal human assistance.
  1. 1.
    We filter the raw caption data from the VG dataset based on metrics such as length. Additionally, we devise an initial prompt that briefly describes the data the LLM (GPT (OpenAI 2023) is used in this paper) should generate with reference to the examples. This prompt, along with the necessary context (caption text), is then used to generate a small sample set. Next, we inspect the generated samples and modify the prompt accordingly. Through several iterations, the refined prompt effectively guides the LLM in generating caption labels. The final prompt is described in Listing 1.
  2. 2.
    Although the labeling data roughly meets expectations, it faces issues due to the limitations of existing LLMs. These issues include confusion between singular and plural forms, overly fine divisions of category labels, and insufficient samples within certain categories. Therefore, the labeled data requires further processing. Initially, we obtain noun lemmas through lemmatization and merge labels derived from the same word. Subsequently, the word2vec model (Mikolov et al. 2013) is employed to derive the word embeddings for the labels, which are then utilized to calculate the similarity between them. Then, manual rules are set, including a similarity threshold and a minimum number of samples within a category, to cluster and filter the labels. By iterating these rules based on the merged data, the final dataset is obtained.
Listing 1 The final prompt used in the MOInst construction
1You will label each sentence of the given multiple sentences divided with <CAP>, the label indicates the object of the sentence. Labeling examples are given.
2
3labeling examples:
4a red car is at the back <CAP> two lights are on on the building <CAP> bikes are parked on the far end <CAP> the word \"walk\" is lit on a street sign <CAP> two men stand on an urban sidewalk
5car <LAB> light <LAB> bicycle <LAB> character <LAB> man
6
7sentences:
8{input_text}
9
10Indicate the subject of each sentence with prototype of one word separately, show me the labels directly without the sentences with each label separated by <LAB>, and make sure that the marked labels correspond one-to-one with the original sentence."
Image ID Image Category ID Category Object ID Captions
484 [Uncaptioned image] 2 fence 4840002 A fence made of metal by the road.
31 bike 4840031 A blue bike is parked off to the side.
36 woman 4840036 A woman is sitting down in a chair.
130 grass 4840130 The grass is green on the side where the woman is sitting.
Table 7: Important information about an image in the MOInst dataset. Image ID, Category ID, and Object ID are unique identifiers for images, category labels, and specific categories within a particular image, respectively.
Experimental Setup
We introduce a new task for which no prior experimental data exists. As a result, we need to evaluate all models, including baselines and our proposed method. However, due to limited computational resources, the cost of running multiple experiments to obtain error bars is prohibitively high. Consequently, it should be noted that the results presented in this paper are based on a single experiment.
Evaluation Metrics
To obtain an accurate measure of visual encoding ability, it is essential to exclude the inflated values introduced by the instructional text. A way to quantify the impact of instructional information on the assessment is to calculate the probability that the results of task execution are consistent with an erroneous instruction. Specifically, for task $\tau(\cdot)$ and model $\Phi(\cdot,\cdot)$, a set of $p$ evaluation data is given as $\{x^{I}_{k},x^{IT}_{k},t_{k}\}^{p}_{k=1}$, where $x^{IT}$ represents the instruction text corresponding to the image $x^{I}$, and $t$ is the ground truth. The evaluation score of the model can be formalized as
$\mathcal{S}=\mathbb{E}[\mathbb{I}(\tau(\Phi(x^{I}_{k},x^{IT}_{k}))=t_{k})].$ (15)
Then the interference of the instruction text on the score can be expressed as
$\mathcal{N}=$ $\mathcal{S}(x^{I}_{k},x^{IT}_{j},t_{j}\mid j\neq k)$ (16)
$-\mathcal{S}(x^{I}_{k},x^{IT}_{j},t_{i}\mid j\neq k,i\in\{1,2,\ldots,p\}),$ (17)
which is the likelihood of being misdirected to a specific wrong minus randomness. Thus, the score obtained since the visual features are
$\mathcal{S}^{V}=\mathcal{S}-\mathcal{N}.$ (18)
Hyper-parameter Settings
We report the hyper-parameter settings used during GiVE training in Table 8. All other settings align with those of the corresponding visual encoder.
Item Value Item Value
Optimizer AdamW Batch size 640
Learning rate 5e-5 Weight decay 3e-3
Epoch 5 Warmup ratio 3e-2
Table 8: Hyper-parameters for training the model.
Compute Resources
We conduct our experiments on a NVIDIA A800-SXM4-80GB GPU, an Intel(R) Xeon(R) Platinum 8358P CPU @ 2.60GHz. The experimental device possesses 1.0 terabyte (TB) of memory, and its operating system is Ubuntu 22.04.3 LTS.
Parameter Sensitivity
Configurations of CLIP with GiVE
We conduct experiments with different configurations of CLIP, and the results in Table 9 demonstrate that models integrated with the AG-Adapter can still be enhanced using typical methods. Specifically, smaller patch sizes can capture more fine-grained information, and an increased number of parameters can provide greater learning capacity. These enhancements improve CLIP’s performance, and similarly, AG-CLIP, which combines CLIP with the GiVE, also benefits from these improvements.
Type PS # TP MOInst LVISf LVISa
I2T T2I F1 AUC F1 AUC
R@1 R@1
B/32 $32^{2}$ 62M 29.1 29.8 50.7 75.6 56.2 77.3
B/16 $16^{2}$ 62M 31.2 33.5 55.8 77.4 61.9 79.9
L/14 $14^{2}$ 240M 34.7 36.9 60.4 80.2 62.8 80.7
Table 9: Evaluation of the GiVE applied on different configurations of CLIP-ViT. “Type” refers to the type of the base model, “PS” indicates the patch size used when dividing the image into patches, and “# TP” represents the number of trainable parameters.
Sensitivity Analysis on Loss Weights
We also explore the impact of different loss weightings within GiVE on the AG-Adapter to balance the richness of effective information and the accuracy of semantic concepts. As observed from results in Table 10, variations in the weightings of these three hyper-parameters cause the visual encoder’s capabilities to favor one aspect of the trade-off over another. Taking everything into account, we adopt the relatively balanced values from the first row.
$\omega^{\text{OITC}}$ $\omega^{\text{OIIC}}$ $\omega^{\text{OID}}$ MOInst LVISf LVISa
I2T T2I F1 AUC F1 AUC
R@1 R@1
1 1 0.01 29.1 29.8 50.7 75.6 56.2 77.3
0.1 1 0.01 15.0 18.4 41.3 71.2 49.9 73.9
0.5 1 0.01 24.7 26.6 51.7 73.1 55.1 76.5
5 1 0.01 30.2 30.4 39.1 68.7 48.9 73.0
10 1 0.01 29.7 30.6 33.2 65.3 42.9 69.5
1 0.1 0.01 29.5 31.0 48.0 73.5 56.6 77.2
1 0.5 0.01 29.3 30.8 48.2 74.1 54.7 76.5
1 5 0.01 24.8 26.7 43.3 71.9 52.5 75.6
1 10 0.01 23.6 27.5 39.4 70.7 49.2 74.3
1 1 0.001 28.9 30.8 32.4 65.1 43.3 69.8
1 1 0.005 29.4 30.0 44.6 71.6 54.3 75.6
1 1 0.05 22.4 24.1 53.3 77.0 54.9 76.6
1 1 0.1 20.4 21.7 48.2 74.2 53.2 75.8
Table 10: Hyper-parameter grid search on $\omega^{\text{OITC}}$, $\omega^{\text{OIIC}}$, and $\omega^{\text{OID}}$, which represent the weights for $\mathcal{L}^{\text{OITC}}$, $\mathcal{L}^{\text{OIIC}}$, and $\mathcal{L}^{\text{OID}}$ , respectively.
Comparison with Clip-based Method
We compare CLIP (Radford et al. 2021), the recent MetaCLIP (Xu et al. 2024), and CLIP with our GiVE applied on five datasets for zero-shot image classification. The results, shown in Table 11, demonstrate that our method enables CLIP to outperform the state-of-the-art (SoTA) method.
Model LVISf LVISa COCO SUN397 ImageNet
AUC AUC AUC AUC AUC
CLIP 57.8 55.5 63.8 81.9 80.2
MetaCLIP 57.8 55.6 63.7 83.6 81.1
CLIP +GiVE 75.6 77.3 75.0 85.2 82.7
Table 11: Top-1 AUC (%) of zero-shot image classification on 5 datasets.
Training Stability
We apply our GiVE method to CLIP-ViT and repeat the experiment 5 times to calculate the mean and standard deviation, as recorded in Table 12. The results demonstrate that our method is relatively stable across all 3 tasks.
Model MOInst LVISf LVISa
I2T T2I AUC AUC
R@1 R@1
CLIP +GiVE $29.1\!\pm\!0.2$ $30.1\!\pm\!0.3$ $75.3\!\pm\!2.2$ $77.4\!\pm\!1.1$
Table 12: The means and standard deviations measured by CLIP with GiVE across 5 experiments on 3 different tasks.
More Guided Attention Visualization
In Figure 7, we use EigenCAM (Muhammad and Yeasin 2020) to generate heat maps that visualize instruction following. The results demonstrate that our proposed method enables existing image encoder to flexibly perceive the query objects.
Figure 7: More visualizations on different categories. Each row represents a set, comprising the raw image and the visualized attention of the image encoders. The first column contains the raw image, the second column shows the attention of CLIP-ViT, and the last four columns illustrate how GiVE guides CLIP-ViT to attend to both salient and non-salient objects.
Visual Attention and Response
Table 13 illustrates the disparate responses of MLLMs when queried with respect to disparate objects. The attention distributions of the visual encoders on images are visualized using EigenCAM (Muhammad and Yeasin 2020), with the redness intensity indicating the degree of attention the network allocates. The table presents a case study where the model accurately responds to a query about a highly salient SUV but incorrectly answers a query about a less noticeable car that shares features with the SUV. This error arises because the model’s perception is constrained by image features that emphasize salient objects. When query information is incomplete, the model tends to default to common sense embedded in its weights. While this approach generally works for abstract text, it can lead to errors when dealing with detail-rich images.
Visual Input Example, Salience and Non-Salience Object Question:
Input image [Uncaptioned image]
CAM
LLaVA-1.5
InstructBLIP
User Describe in detail the SUV on the right.
LLaVA-1.5 (Liu et al. 2023a)
The SUV on the right is a yellow vehicle, possibly a van or a truck, with a man standing on the back of it. The man is holding a shirt, possibly ironing it or preparing to hang it on a clothesline. The SUV is parked on the side of the road, and it appears to be a unique and unconventional sight.
InstructBLIP (Dai et al. 2023)
In the image, a yellow SUV is parked on the side of a busy city street. A man wearing a yellow shirt is standing on the back of the SUV, ironing clothes on an ironing board that is attached to the vehicle’s roof rack. He is positioned in such a way that he can easily access the ironing board from the back of the SUV. The presence of the ironing board on top of the SUV adds a unique and creative touch to the scene, showcasing the man’s resourcefulness and his ability to make the most out of his surroundings. Additionally, two taxis are visible in the background, adding to the bustling atmosphere of the city street.
User Describe in detail the car on the left.
LLaVA-1.5 (Liu et al. 2023a)
The car on the left is a yellow taxi cab. It is driving down the street, and there is a man standing on the back of the car, holding a shirt. The taxi is passing by a building, and the man appears to be hanging the shirt on a clothesline.
InstructBLIP (Dai et al. 2023)
In the image, a man wearing a yellow shirt is seen standing on top of a car parked on the side of a busy city street. The car appears to be a large SUV or minivan, possibly a Toyota Highlander or Honda Odyssey. The man is ironing clothes on the roof of the vehicle, which is an unusual and creative way to do laundry in the middle of a bustling city. The presence of a yellow taxi cab nearby adds to the vibrant and lively atmosphere of the scene.
Table 13: Responses of MLLMs to queries on objects with varying attention levels.
In comparison, our GiVE aids in addressing the visual overlook issue. It enables the visual encoder to focus on the query object, even if it is non-salient. We substitute AG-CLIP-ViT-L/14, as referenced in Table 9, for CLIP-ViT-L/14, which is utilized by LLaVA (Liu et al. 2023b), without additional training. As illustrated in Table 14, this methodology allows LLaVA to generate a more accurate response.
Visual input example, Non-Salience Object Question:
User Describe in detail the car on the left.
LLaVA w/o GiVE The car on the left is a yellow taxi cab. It is driving down the street, and a man is standing on the back of the car, holding a clothes iron. This unusual scene suggests that the man might be a tailor or a professional ironing service provider, offering his services to the people in the city.
LLaVA w/ GiVE The car on the left is a yellow taxi cab. It is driving down the street, passing by a man who is ironing clothes in the back of a truck. The taxi cab is positioned behind the truck, and it appears to be in motion.
Table 14: The integration of GiVE enables MLLM to utilize more targeted visual features, thereby enhancing response quality. In the example, LLaVA with GiVE does not incorrectly position the man behind the car, and the response focuses on the car rather than mistakenly shifting emphasis to the man, as seen with the original LLaVA.

### v2

@@ II-D Instruction of New Dataset
AG-CLIP 62M CLIP-ViT B/32 29.1 29.8 50.7 75.6 56.2 77.3
III Experiments
III-A Experimental Setup
III-A1 Datasets
We evaluate the effectiveness of GiVE using both the LVIS dataset [15] and MOInst dataset, each annotating multiple objects per image. The comprehensive LVIS dataset, referred to as LVISa, comprises 1,203 categories, from which, we derive a subset, denoted as LVISf, with 405 “frequent” categories. In contrast, the MOInst dataset used for training contains 264 categories, with category label overlaps of 7.7% and 17.6% with the two LVIS datasets, respectively. We also conduct experiments on the COCO [16], SUN397 [17], and ImageNet [18] datasets.
III-A2 Baselines
We evaluate the gains brought by integrating the GiVE with several representative ViT baselines, including CLIP [10], GroupViT [13], SigLIP [19], OwlViT [20], and MetaCLIP [21]. We also apply the GiVE to larger encoding frameworks, such as BLIP-2 [1] and InstructBLIP [22]. Throughout this paper, unless explicitly stated otherwise, the experiments with GiVE are conducted on the CLIP platform, with the unmarked CLIP model referring to CLIP-ViT-B/32.
III-A3 Evaluation Metrics
Following previous research [23], we evaluate image classification and image-text retrieval tasks using scores based solely on vector similarity, avoiding further optimization to accurately reflect the extracted feature information. To focus on the capabilities of visual encoders, we quantify and mitigate the influence of textual data in methods like InstructBLIP [22].
TABLE IV: Ablation study
Loss Fusion11footnotemark: 1 Inst. MOInst LVISf LVISa
@@ III-A3 Evaluation Metrics
-OITC22footnotemark: 2 Early33footnotemark: 3+late44footnotemark: 4 0.1 0.1 0.3 50.0 0.1 50.0
@@
All Early+late (dense66footnotemark: 6) 29.1 29.8 50.7 75.6 56.2 77.3
  • 1
    The layer where the AG-Adapter is inserted.
  • 2
    “-” indicates the elimination of the loss function during training.
  • 3,4,5,6
    “Early”, “late”, “sparse”, and “dense” indicate that features are fused in the first half of layers, the latter half of layers, alternate layers, and all layers, respectively.
III-B Performance Evaluation
III-B1 Image Classification
We conduct zero-shot image classification on the “frequent” subset and the full LVIS test set. Table I presents the results on five ViT baselines. Key observations from these experiments are as follows:
@@ III-B1 Image Classification
III-B2 Image-Text Retrieval
Image-text retrieval includes two subtasks: image-to-text retrieval and text-to-image retrieval. We evaluate the models on the MOInst dataset. Since the AG-Adapter is trained on this dataset, we compare the results under the fine-tuned setting, as recorded in Table II. It is easy to see that GiVE outperforms all baseline methods while utilizing fewer trainable parameters. The informativeness gap between instruction and caption indicates that simple instruction alone is insufficient for the model to achieve such a high hit rate. Together, GiVE effectively extracts target visual features. Additionally, the evaluation results for “Instruct.” further support the credibility of our experimental findings.
III-B3 Comparison with Secondary Coding Methods
The BLIP-2 and InstructBLIP methods utilize a Q-Former to re-encode image features after the visual encoder, with InstructBLIP incorporating text instructions into the Q-Former. We replace the visual encoders of these two methods with AG-CLIP-ViT-B/32 and then fine-tune them on MOInst following their respective strategies. The results are shown in Table III, where “Inst.” indicates whether the model receives additional instructions. “Img Enc.” specifies the image encoder used by the model, and “Type” denotes the type of image encoder. “GiVE” shows whether our GiVE method is applied to the image encoder, and “Post Proc.” identifies the type of post-processor used for the secondary encoding of the image embeddings. From the table, it can be observed that:
@@ III-B3 Comparison with Secondary Coding Methods
III-C Ablation Studies
We perform an ablation study of GiVE from three aspects: loss, fusion layers, and abstract semantic instructions. For instructions, we replace short object prompts with detailed descriptions to remove abstract semantics. Table IV shows the results, with the gray row for CLIP-ViT and the last row for AG-CLIP-ViT. The main observations are as follows:
@@ III-C Ablation Studies
IV Conclusion